# references.bib

### v1

@@ Abstract
As the burden of herbicide resistance grows and the environmental repercussions of excessive herbicide use become clear, new ways of managing weed populations are needed. This is particularly true for cereal crops, like wheat and barley, that are staple food crops and occupy a globally significant portion of agricultural land. Even small improvements in weed management practices across these major food crops worldwide would yield considerable benefits for both the environment and global food security. Blackgrass is a major grass weed which causes particular problems in cereal crops in north-west Europe, a major cereal production area, because it has high levels of of herbicide resistance and is well adapted to agronomic practice in this region. With the use of machine vision and multispectral imaging, we investigate the effectiveness of state-of-the-art methods to identify blackgrass in wheat and barley crops. As part of this work, we provide a large dataset with which we evaluate several key aspects of blackgrass weed recognition. Firstly, we determine the performance of different CNN and transformer-based architectures on images from unseen fields. Secondly, we demonstrate the role that different spectral bands have on the performance of weed classification. Lastly, we evaluate the role of dataset size in classification performance for each of the models trialled. We find that even with a fairly modest quantity of training data an accuracy of almost 90% can be achieved on images from unseen fields.
@@ 1 Introduction
Weeds compete with crops for light, water and nutrients. Additionally, grains from grass weeds get harvested alongside grains from the crop, contaminating the harvest, and can drastically lower the price received for the crop. Current mainstream weed management practices involve broadcast spraying, where the entire field is sprayed with a selective herbicide. An alternative approach, known as precision spraying, targets herbicide only at areas of weed cover guided by machine vision. Precision spraying facilitates up to 95% reduction in herbicide use (depending on crop and field) [zanin2022reduction], and a proportional decrease in herbicide costs and environmental damage from runoff. In addition, precision spraying allows a wider range of herbicides (potentially broad spectrum) to be used, opening more options for resistance management [busi2020rotations]. Meanwhile, other technologies, like automated camera-guided mechanical weeding, promise non-chemical weed control. Both of these promising technologies are predicated on accurate, real-time weed recognition within often complex crop canopies.
@@
Efforts to reduce herbicide usage in staple cereal crops have the potential to deliver significant impact. The area of global cropland devoted to growing cereals, and in particular rice, wheat and maize, is orders of magnitude greater than that of many vegetable crops [fao]. Furthermore, wheat is second only to rice as a global staple [Eren2022], with a global consumption of 65.6 kg per person per year. Therefore, not only is sustainable and reliable crop care for wheat critical for global food security, any technology that reduces herbicide usage in these global staples has the potential to massively reduce the environmental impact of herbicide use globally.
@@
Moreover, particular weeds present a greater threat to global food supply than others; grass weeds are a particular problem in wheat production due to their biological similarities [chhokar2012weed, young1996weed, moyer1994weed, Hick2018]. In Europe, blackgrass is one of the most economically damaging weeds [Vara2020], so effective strategies to manage populations are a priority. Due to high levels of resistance to selective herbicides [Hick2018], using them for precision spraying may not be sufficient to control blackgrass populations. In these instances, precision spraying with non-selective herbicides could be an alternative approach, but requires a high degree of confidence in detection and targeting so as not to damage the crop. Such confidence will rely on well tested, accurate, real time machine vision systems that can be deployed at scale to the field.
@@ 2.1 Color Spaces for Weed Recognition
Initially, color indices derived from RGB like Excess Greenness (ExG), Excess Redness (ExR), ExG minus ExR (ExGR) and Normalized Difference Index (NDI) were used in weed recognition research [jensen2021predictive]. Multispectral imaging for crop and weed discrimination is now an active area of research. There is an existing precedent for the use of multispectral imaging in plant analysis since the Normalized Difference Vegetation Index (NDVI) [rouse1974monitoring], derived from near-infrafed (NIR) and red light, has long been used in agriculture to assess crop health. Studies trialling different combinations of NIR, Red and NDVI as inputs to a convolutional neural network (CNN) to detect weeds in sugar beet fields found that NIR and Red values used together performed the best when identifying weeds [sa2018weednet, razaak2019integrated]. Interestingly, including NDVI as an input, alongside NIR and Red values, did not improve the performance [sa2018weednet, razaak2019integrated]. With the use of hyperspectral cameras, it was found that increasing the number of bands from 3 up to 61 resulted in incremental improvements in weed species classification accuracy [farooq2018analysis]. In a study where several color indices where trialled, the inclusion of NIR information either alone, or alongside other indices like ExG and NDI, significantly improved the segmentation accuracy [wang2020semantic].
@@ 2.2 Weed Recognition Techniques
Early research into techniques for weed recognition relied on hand-crafted feature-based techniques. Approaches to discriminating grass weeds in cereal crops use color [anderegg2023farm], texture [xu2020recognition], depth [xu2020recognition] and spectral [amziane2021weed, jensen2021predictive] features. For weed classification, support vector machine (SVM) [cristianini2000introduction] and k-Nearest Neighbor (k-NN) [cover1967nearest] approaches were used to learn decision boundaries in the feature space [xu2020recognition].
@@
Deep learning-based weed classification rose to prominence after the success of early CNN architectures like AlexNet [krizhevsky2012imagenet] and VGG [simonyan2014very]. However, ResNet [he2016deep] improved the applicability of this technology by utilising skip connections to address the vanishing gradient problem, which enabled the training of deeper neural networks, resulting in improved performance and convergence. In the following years, deep learning-based methods with faster inference speeds and smaller memory footprints enabled weed recognition on the edge, like the popular MobileNetV3 architecture, which exploits lightweight depthwise separable convolutions for improved efficiency [howard2019searching]. EfficientNet employed a neural architecture search approach to design a network and a novel method to scale the model depth, width, and resolution of the network in order to balance accuracy and efficiency[tan2019efficientnet]. Recently, transformers have demonstrated how powerful self-attention mechanisms can be for recognizing complex patterns in sequential data [vaswani2017attention]. Vision Transformer (ViT) [dosovitskiy2020image] applied transformer technology to image classification by treating an image as a sequence of patches, which were then linearly embedded and processed through a transformer architecture. The self-attention mechanism in transformers allowed ViT to analyze global contextual information, improving performance on various image classification benchmarks.
@@
ResNet, MobileNetV3 and MobileViT [mehta2021mobilevit] — a compact version of ViT — demonstrated strong performance in weed classification in wheat [alirezazadeh2023comparative]. Deep learning-based object detection techniques, that identify the region where an object is located in an image, have been used to detect grass weeds in wheat crops including using DetectNet [dyrmann2017roboweedsupport], SSD [dyrmann2018using], Faster R-CNN [xu2021multi, xu2023weedsnet]. Additionally, there is work segmenting grass species, such as ryegrass, using an adapted ERFNet [su2021real], as well as green foxtail and horsetail using an adapted U-Net network [zou2021modified].
@@ 2.3 Prevalence of Grass Weed Recognition in Cereal Crops
We systematically reviewed the latest machine vision research on wheat and barley crops and identified that grass weeds are understudied compared to broadleaf weeds. Grass weeds are more challenging to distinguish from cereal crops (which are themselves grass species). Papers were identified in Google Scholar and Scopus using search query “weed identification” OR “weed recognition” OR “weed discrimination” OR “weed detection” OR “weed classification” OR “weed segmentation” on 20th October 2023. All publications from recognised, peer-reviewed, English-language journals and conference venues that could be accessed were included. For this analysis, we are only interested in weed detection in wheat and barley crops, so excluding all other weed detection papers we found a total of 79 papers.
@@
Figure 1 shows that broadleaf weeds have garnered more research interest than grass weeds. One possible explanation for this discrepancy is how challenging weed recognition is in grass weeds compared to broadleaf weeds. Grass crops also present a high level of challenge compared to broad leaf crops as they occlude one another, as well as any weeds in the field, even at relatively early growth stages. By contrast, many broadleaf crops remain spaced out, display distinctive features, for example distinctive color and leaf shape, at early growth stages. Additionally, many grass crops look very similar to other grass plants. Distinguishing between grass species is an example of fine-grained visual classification [ma2021research]. This is a subgroup of classification tasks that aims to distinguish between similar examples based on some subcategory like plant species or car model. Moreover, like with the classification of medical images, another fine-grained visual classification task, experts are required to produce ground-truth images — adding to the cost of producing a dataset. Therefore, grass weed recognition in cereal crops has some of the same challenges as medical imaging domains when it comes to building datasets.
Figure 1: Weed detection research effort in wheat and barley across weed Species, as measured by published papers, for broad leaf (blue) and grass (red) weeds.
@@
Figure 2: Example images from the Plant Seedlings Dataset [giselsson2017public].
An early contribution to the field of deep learning for weed detection in wheat was the Plant Seedling Classification dataset [giselsson2017public]. This dataset contained images of crops including wheat, maize and sugar beet along with 9 common weed species including blackgrass. Over a third of the papers on wheat/weed recognition reviewed for this survey carried out experiments on the Plant Seedling Classification dataset [chavan2018agroavnet, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet]. This dataset has made a great contribution to the field by drawing attention to the problem of crop and weed discrimination for crops that are critical to the world’s food supply. However, using this dataset does not enable an evaluation of in-field crop and weed discrimination.
The rest of the wheat papers use in-field images of wheat and weeds. A few papers explore the problem of weed detection in mature wheat crops using both color features [rasmussen2019pre] and LiDAR [shahbazi2021assessing]. Most of the others attempt to detect weeds at an early stage. Many of these focused solely on broadleaf weeds [pflanz2018weed, alirezazadeh2023comparative, el2022metaheuristic, zhuang2022evaluation, almalky2022efficient, almalky2023deep, almalky2023real, xu2023weedsnet] which are easier to distinguish in grass crops than grass weeds. Some of the work discriminating grasses, hand-crafted approaches using color [anderegg2023farm], texture [xu2020recognition], depth [xu2020recognition] and spectral [amziane2021weed, jensen2021predictive] features were implemented. In this work, we focus on using deep learning-based approaches as they have been shown to be more robust to changes in illumination and occlusion.
2.4 Paper Overview
To facilitate evaluating agricultural vision systems on in-context, in-field data, we present a weed image data set featuring wheat and barley of over 15,000 images across 51 fields in the UK, with and without blackgrass — a grass weed of cereal crops. This dataset is comparable in scale to the seminal Plant Seedling Classification dataset [giselsson2017public]. The visual similarity of blackgrass to cereal crops poses a challenging fine-grained classification task for precision weed management technology. The release of this dataset will accelerate the development of weed detection systems for black-grass, and due to their visual similarity, other grass weeds, for wheat, a critical global staple.
The contributions of this paper are as follows:
  • A dataset containing over 15,000 multispectral images of wheat and blackgrass stratified by field, season, geo-location and soil-type, enabling the study of domain generalisation with respect to these factors;
@@ 3 Materials and Methods
In this section, we first present our dataset of blackgrass in wheat and barley crops. We then outline our detection approach including architectures, experiments, training settings and evaluation methods.
@@ 3.1 Dataset
The dataset [^1] The dataset will be made publicly available upon acceptance. consists of images of blackgrass in wheat and barley crops collected from 51 fields across eight different soil types in the east of England, a key grain growing region of the UK, between 3rd November 2020 and 16th October 2021. An average of 300 images was taken from each field. The majority of fields had more than 100 images both with and without blackgrass, however, seven fields only had ‘blackgrass’ images and four fields only had ‘no blackgrass’ images. As shown in Table 1, the images were split into by field so that there is approximately a 80-10-10 split for training, validation and test sets. The aim of this stratification is to ensure that each model’s ability to generalise to unseen fields is assessed. Table 3.1 shows how many images are in each set according class, crop type and season. Care was taken to ensure the sets were balanced in terms of positive and negative examples as well as to ensure each set contain some images from each stage in the season and each crop type.
@@
{tblr}
—c—c—c—c—c—c— Train Validation Test Total
[…]
@@
[…]
@@
Land managers were consulted to find patches of blackgrass in wheat and barley fields. Images of class ‘no blackgrass’ were taken near the blackgrass patches. Fields with a good number of blackgrass patches were re-visited to collect images at different growth stages of the crop and blackgrass. Blackgrass class was assessed at the time of image collection by the same observer for all images across all fields to maximise the number of images collected. The dataset therefore only has image level labels.
@@
Figure 3: RGB examples from the dataset. In a) blackgrass are the thin single green strands and the wheat is the wider leaf blades. Blue arrows show a few indicative examples of blackgrass, but there are many blackgrass seedlings in a)
For each field, the soil type in that field was categorised according to the Cranfield Soil and Agrifood Institute’s Soilscapes databases [soilscapes]. Due to the geography of the farms that were visited, there were no fields visited that contained predominantly peat, or predominantly chalky soil. Table 3 describes each soil type in the dataset and shows the break down of images with each soil type in each of the dataset splits. Note that, due to prioritising a stratification based on geo-location, it was not possible to have every soil type represented in each dataset split.
@@
Type 1 Free draining lime-rich loamy 379 0 0 379
Type 2 Free draining, base-rich, slightly acidic loamy 497 0 581 1078
Type 3 Wet slightly acidic loamy and clayey 489 0 492 981
Type 4 Wet lime-rich loamy and clayey 1724 0 0 1724
Type 5 Acidic sandy and loamy 0 0 0 0
Type 6 Base-rich loamy and clayey, wet 1444 214 0 1658
Type 7 Loamy and clayey wet 6204 1101 197 7502
Type 8 Wet loamy and sandy, peaty surface 1015 744 848 2607

### v2

# Multispectral Fine-Grained Classification of Blackgrass in Wheat and Barley Crops

@@ Abstract
As the burden of herbicide resistance grows and the environmental repercussions of excessive herbicide use become clear, new ways of managing weed populations are needed. This is particularly true for cereal crops, like wheat and barley, that are staple food crops and occupy a globally significant portion of agricultural land. Even small improvements in weed management practices across these major food crops worldwide would yield considerable benefits for both the environment and global food security. Blackgrass is a major grass weed which causes particular problems in cereal crops in north-west Europe, a major cereal production area, because it has high levels of herbicide resistance. Moreover, detecting blackgrass in grass crops is challenging due to its visual similarity to these crops. Despite this, a systematic review of the literature on weed recognition in wheat and barley, included in this study, highlights that blackgrass – and grass weeds more broadly – have received less research attention compared to certain broadleaf weeds. With the use of machine vision and multispectral imaging, we investigate the effectiveness of state-of-the-art methods to identify blackgrass in wheat and barley crops. As part of this work, we present the Eastern England Blackgrass Dataset, a large dataset with which we evaluate several key aspects of blackgrass weed recognition. Firstly, we determine the performance of different CNN and transformer-based architectures on images from unseen fields. Secondly, we demonstrate the role that different spectral bands have on the performance of weed classification. Lastly, we evaluate the role of dataset size in classification performance for each of the models trialled. All models tested achieved an accuracy greater than 80%. Our best model achieved 89.6% and that only half the training data was required to achieve this performance. Our dataset is available at: https://lcas.lincoln.ac.uk/wp/research/data-sets-software/eastern-england-blackgrass-dataset/.
@@ 1 Introduction
Weeds compete with crops for light, water and nutrients. Additionally, grains from grass weeds get harvested alongside grains from the crop, contaminating the harvest, and can drastically lower the price received for the crop. Current mainstream weed management practices involve broadcast spraying, where the entire field is sprayed with a selective herbicide. An alternative approach, known as precision spraying, targets herbicide only at areas of weed cover guided by machine vision. Precision spraying facilitates up to 95% reduction in herbicide use (depending on crop and field) zanin2022reduction , and a proportional decrease in herbicide costs and environmental damage from runoff. In addition, precision spraying allows a wider range of herbicides (potentially broad spectrum) to be used, opening more options for resistance management busi2020rotations . Meanwhile, other technologies, like automated camera-guided mechanical weeding, promise non-chemical weed control. Both of these promising technologies are predicated on accurate, real-time weed recognition within often complex crop canopies.
@@
Efforts to reduce herbicide usage in staple cereal crops have the potential to deliver significant impact. The area of global cropland devoted to growing cereals, and in particular rice, wheat and maize, is orders of magnitude greater than that of many vegetable crops fao . Furthermore, wheat is second only to rice as a global staple Eren2022 , with a global consumption of 65.6 kg per person per year. Therefore, not only is sustainable and reliable crop care for wheat critical for global food security, any technology that reduces herbicide usage in these global staples has the potential to massively reduce the environmental impact of herbicide use globally darbyshire2024review .
@@
Moreover, particular weeds present a greater threat to global food supply than others; grass weeds are a particular problem in wheat production due to their biological similarities chhokar2012weed ; young1996weed ; moyer1994weed ; Hick2018 . In Europe, blackgrass is one of the most economically damaging weeds Vara2020 , so effective strategies to manage populations are a priority. Due to high levels of resistance to selective herbicides Hick2018 , using them for precision spraying may not be sufficient to control blackgrass populations. In these instances, precision spraying with non-selective herbicides could be an alternative approach, but requires a high degree of confidence in identifying and targeting so as not to damage the crop. Such confidence will rely on well tested, accurate, real time machine vision systems that can be deployed at scale to the field.
1.1 Paper Overview
To facilitate evaluating agricultural vision systems on in-context, in-field data, we present the Eastern England Blackgrass Dataset, a weed image data set featuring wheat and barley of over 15,000 images across 51 fields in the UK, with and without blackgrass — a grass weed of cereal crops.
Unlike other crop and weed datasets, this dataset focuses on recognising the presence of blackgrass in wheat. The visual similarity of blackgrass to cereal crops poses a challenging fine-grained classification task for precision weed management technology. The release of this dataset will accelerate the development of weed identification systems for blackgrass, and due to their visual similarity, other grass weeds, for wheat, a critical global staple.
The contributions of this paper are as follows:
  • A dataset containing over 15,000 multispectral images of wheat and blackgrass stratified by field, season, geolocation and soiltype, enabling the study of domain generalisation with respect to these factors;
@@ 2.1 Color Spaces for Weed Recognition
Initially, color indices derived from RGB like Excess Greenness (ExG), Excess Redness (ExR), ExG minus ExR (ExGR) and Normalized Difference Index (NDI) were used in weed recognition research jensen2021predictive . Multispectral imaging for crop and weed discrimination is now an active area of research. There is an existing precedent for the use of multispectral imaging in plant analysis since the Normalized Difference Vegetation Index (NDVI) rouse1974monitoring , derived from near-infrafed (NIR) and red light, has long been used in agriculture to assess crop health. Studies trialling different combinations of NIR, Red and NDVI as inputs to a convolutional neural network (CNN) to detect weeds in sugar beet fields found that NIR and Red values used together performed the best when identifying weeds sa2018weednet ; razaak2019integrated . Interestingly, including NDVI as an input, alongside NIR and Red values, did not improve the performance sa2018weednet ; razaak2019integrated . With the use of hyperspectral cameras, it was found that increasing the number of bands from 3 up to 61 resulted in incremental improvements in weed species classification accuracy farooq2018analysis . In a study where several color indices where trialled, the inclusion of NIR information either alone, or alongside other indices like ExG and NDI, significantly improved the segmentation accuracy wang2020semantic .
@@ 2.2 Weed Recognition Techniques
Early research into techniques for weed recognition relied on hand-crafted feature-based techniques. Approaches to discriminating grass weeds in cereal crops use color anderegg2023farm , texture xu2020recognition , depth xu2020recognition and spectral amziane2021weed ; jensen2021predictive features. For weed classification, support vector machine (SVM) cristianini2000introduction and k-Nearest Neighbor (k-NN) cover1967nearest approaches were used to learn decision boundaries in the feature space xu2020recognition .
@@
Deep learning-based weed classification rose to prominence after the success of early CNN architectures like AlexNet krizhevsky2012imagenet and VGG simonyan2014very . However, ResNet he2016deep improved the applicability of this technology by utilising skip connections to address the vanishing gradient problem, which enabled the training of deeper neural networks, resulting in improved performance and convergence. In the following years, deep learning-based methods with faster inference speeds and smaller memory footprints enabled weed recognition on the edge, like the popular MobileNetV3 architecture, which exploits lightweight depthwise separable convolutions for improved efficiency howard2019searching . EfficientNet employed a neural architecture search approach to design a network and a novel method to scale the model depth, width, and resolution of the network in order to balance accuracy and efficiencytan2019efficientnet . Recently, transformers have demonstrated how powerful self-attention mechanisms can be for recognizing complex patterns in sequential data vaswani2017attention . Vision Transformer (ViT) dosovitskiy2020image applied transformer technology to image classification by treating an image as a sequence of patches, which were then linearly embedded and processed through a transformer architecture. The self-attention mechanism in transformers allowed ViT to analyze global contextual information, improving performance on various image classification benchmarks.
@@
ResNet, MobileNetV3 and MobileViT mehta2021mobilevit — a compact version of ViT — demonstrated strong performance in weed classification in wheat alirezazadeh2023comparative . Deep learning-based object detection techniques, that identify the region where an object is located in an image, have been used to detect grass weeds in wheat crops including using DetectNet dyrmann2017roboweedsupport , SSD dyrmann2018using , Faster R-CNN xu2021multi ; xu2023weedsnet . Additionally, there is work segmenting grass species, such as ryegrass, using an adapted ERFNet su2021real , as well as green foxtail and horsetail using an adapted U-Net network zou2021modified .
2.3 Weed Recognition Datasets
In recent years, there have been several popular crop and weed datasets. Some have image-level labels such as the Plant Seedlings Dataset giselsson2017public , some have bounding box labels like the Lincoln Beet dataset salazar2022beyond and some have pixel-level labels like the CWFID dataset haug2015crop . As the granularity of the labels has increased, from image-level to pixel-level, it has increased the precision with which that weed can be identified and thus treated. In this landscape, we see the problem of identifying grass weeds in cereal crops is generally overlooked. This is despite the identificiation of grass weeds in cereal crops often being more challenging than other weed detection problems, due to significant occlusion, even during early growth stages, and the visual similarity between the crop and weed.
@@ 2.4 Prevalence of Grass Weed Recognition in Cereal Crops
We systematically reviewed the latest machine vision research on wheat and barley crops from 2015 and identified that grass weeds are understudied compared to broadleaf weeds. Grass weeds are more challenging to distinguish from cereal crops (which are themselves grass species). Papers were identified in Google Scholar and Scopus using search query “weed identification” OR “weed recognition” OR “weed discrimination” OR “weed detection” OR “weed classification” OR “weed segmentation” on 20th October 2023. All publications from recognised, peer-reviewed, English-language journals and conference venues that could be accessed were included. For this analysis, we are only interested in weed recognition in wheat and barley crops, so excluding all other weed recognition papers we found a total of 79 papers.
@@
Figure 1 shows that broadleaf weeds have garnered more research interest than grass weeds. One possible explanation for this discrepancy is how challenging weed recognition is in grass weeds compared to broadleaf weeds. Grass crops also present a high level of challenge compared to broad leaf crops as they occlude one another, as well as any weeds in the field, even at relatively early growth stages. By contrast, many broadleaf crops remain spaced out, display distinctive features, for example distinctive color and leaf shape, at early growth stages. Additionally, many grass crops look very similar to other grass plants. Distinguishing between grass species is an example of fine-grained visual classification. This is a subgroup of classification tasks that aims to distinguish between similar examples based on some subcategory like plant species or car model ma2021research . Moreover, like with the classification of medical images, another fine-grained visual classification task, experts are required to produce ground-truth images — adding to the cost of producing a dataset. Therefore, grass weed recognition in cereal crops has some of the same challenges as medical imaging domains when it comes to building datasets.
Figure 1: Weed recognition research effort in wheat and barley across weed Species, as measured by published papers, for broad leaf (blue) and grass (red) weeds.
@@
Figure 2: Example images from the Plant Seedlings Dataset giselsson2017public .
An early contribution to the field of deep learning for weed recognition in wheat was the Plant Seedling Classification dataset giselsson2017public , shown in Figure 2. This dataset contained images of crops including wheat, maize and sugar beet along with 9 common weed species including blackgrass. Over a third of the papers on wheat/weed recognition reviewed for this survey carried out experiments on the Plant Seedling Classification dataset chavan2018agroavnet ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; mishra2023comparative ; mu2023densenet . This dataset has made a great contribution to the field by drawing attention to the problem of crop and weed discrimination for crops that are critical to the world’s food supply. However, using this dataset does not enable an evaluation of in-field crop and weed discrimination.
Apart from the papers using the Plant Seedling Dataset listed above, the rest of the research papers we reviewed (listed in Appendix A) used in-field data. A few papers explore the problem of weed detection in mature wheat crops using both color features rasmussen2019pre and LiDAR shahbazi2021assessing . Most of the others attempt to detect weeds at an early stage. Many of these focused solely on broadleaf weeds pflanz2018weed ; alirezazadeh2023comparative ; el2022metaheuristic ; zhuang2022evaluation ; almalky2022efficient ; almalky2023deep ; almalky2023real ; xu2023weedsnet which are easier to distinguish in grass crops than grass weeds. Some of the work discriminating grasses, hand-crafted approaches using color anderegg2023farm , texture xu2020recognition , depth xu2020recognition and spectral amziane2021weed ; jensen2021predictive features were implemented. In this work, we focus on using deep learning-based approaches as they have been shown to be more robust to changes in illumination and occlusion.
@@ 3 Materials and Methods
In this section, we first present our dataset of blackgrass in wheat and barley crops. We then outline our classification approach including architectures, experiments, training settings and evaluation methods.
@@ 3.1 Dataset
Our dataset [^1] The Eastern England Blackgrass Dataset is available here: https://lcas.lincoln.ac.uk/wp/research/data-sets-software/eastern-england-blackgrass-dataset/, the Eastern England Blackgrass Dataset, consists of images of blackgrass in wheat and barley crops collected from 51 fields across eight different soil types in the east of England, a key grain growing region of the UK, between 3rd November 2020 and 16th October 2021.
Fields were found by identifying and contacting land managers across eastern England. We started with those we have worked with previously, and identified new land mangers by introduction from those land managers we were already involved. Land managers were asked if we could access fields that were growing wheat or barley, and contained areas that had both low and high black grass densities. We worked with land managers to find fields across different soil types.
An average of 300 images was taken from each field. The majority of fields had more than 100 images both with and without blackgrass, however, seven fields only had ‘blackgrass’ images and four fields only had ‘no blackgrass’ images.
Images were captured using Micasense RedEdge-MX. The camera was mounted on an image collection rig kept at 95cm throughout data collection. In addition to RGB, the camera captures two additional bands: Red Edge (717 nm) and near-infrared (NIR) (842 nm). Red Edge and NIR wavelengths were included because they are known to aid in plant species differentiation, as variations in leaf structure influence how these wavelengths are reflected lang2015near .
As shown in Table 1, the images were split into training, test and validation sets by field so that there is approximately a 80-10-10 split for training, validation and test sets. The aim of this stratification is to ensure that each model’s ability to generalise to unseen fields is assessed. Table 2 shows how many images are in each set according class, crop type and season. Care was taken to ensure the sets were balanced in terms of positive and negative examples as well as to ensure each set contain some images from each stage in the season and each crop type.
@@
Train Validation Test Total
Class Blackgrass 5770 1023 1095 7888
@@
Crop Type Winter Wheat 9425 1731 1569 12725
dashed]2-6 Triticale 172 0 0 172
@@
Land managers were consulted to find patches of blackgrass in wheat and barley fields. Images of class ‘no blackgrass’ were taken near the blackgrass patches. Fields with a good number of blackgrass patches were re-visited to collect images at different growth stages of the crop and blackgrass. The assessment of “Blackgrass” or “No Blackgrass” was made at the time of image collection by the same observer for all images across all fields to maximise the number of images collected.
The observer often needed to carefully examine suspected plants for distinguishing features, such as the absence of auricles and jagged ligules. Since these details are not clearly visible in the images, even an expert observer would struggle to re-identify blackgrass from the photos alone. As a result, the dataset contains only image-level labels.
@@
Figure 3: RGB examples from the dataset. In a) blackgrass are the thin single green strands and the wheat is the wider leaf blades. Blue arrows show a few indicative examples of blackgrass, but there are many blackgrass seedlings in a). In c), the blackgrass it the plant towards the top right, with thinner blades, while the wheat plant is towards the lower left with wider blades. b) and c) show examples without blackgrass.
For each field, the soil type in that field was categorised according to the Cranfield Soil and Agrifood Institute’s Soilscapes databases soilscapes . Due to the geography of the farms that were visited, there were no fields visited that contained predominantly peat, or predominantly chalky soil. Table 3 describes each soil type in the dataset and shows the break down of images with each soil type in each of the dataset splits. Note that, due to prioritising a stratification based on geo-location, it was not possible to have every soil type represented in each dataset split.
@@
Type 5 Free draining lime-rich loamy 379 0 0 379
Type 7 Freely draining slightly acid but base-rich soils 497 0 581 1078
Type 8 Slightly acid loamy and clayey soils with impeded drainage 489 0 492 981
Type 9 Lime-rich loamy and clayey soils with impeded drainage 1724 0 0 1724
Type 17 Slowly permeable seasonally wet acid loamy and clayey soils 6204 1101 197 7502
Type 18 Slowly permeable seasonally wet slightly acid but base-rich loamy and clayey soils 1444 214 0 1658
Type 23 Loamy and sandy soils with naturally high groundwater and a peaty surface 1015 744 848 2607